1.) Misu Sheet owner of the bedspread shop knows his customers will pay no more than $100 for a comforter. Misu wants a 40% markup on selling price. What is the most Misu can pay for a comforter?

2.) Angies bake shop makes birthday chocolate chip cookies that cost $2 each. Angie expects that 10% of the cookies will crack and be discarded. Angie wants a 65% markup on cost and produces 90 cookies. What should angie price each cookie?
Answer #1

Selling price = $100

Selling price = Cost price + Mark up

100 = Cost + 40%*Selling price

100 = Cost + 40%*100

Hence, Maximum cost = $60

2)Total cost incurred = 2*90 = $180

Number of good units = 90*90% = 81 cookies

Desired Revenue = 180 + 65%*180 = $297

Hence, price per cookie = 297/81

= $3.6667

i.e. $3.67 per cookie